OpenStudy (anonymous):

Find the horizontal asymptotes of √3x^2+4/5x+3

OpenStudy (anonymous):

does the equation look like this \[\sqrt{3x ^{2}}\div5x +3\]

OpenStudy (anonymous):

\[\lim_{x\to\infty}\frac{\sqrt{3x^2}}{5x+3}=\lim_{x\to\infty}\frac{\sqrt{x^2}\sqrt3}{5x+3}\] \(\sqrt{x^2}=|x|=x\), since \(x\to\infty\) (x is positive). \[\lim_{x\to\infty}\frac{\sqrt3 x}{5x+3}=\lim_{x\to\infty}\frac{\sqrt3}{5+\frac{3}{x}}=\cdots\] Do something similar for \(x\to-\infty\), with the exception that now \(\sqrt{x^2}=|x|=-x\), since x is negative.